OpenStudy (anonymous):

A long distance trucker traveled a 180 miles during a snow storm in one direction. The return trip on a rainy day at double the speed it took 3 hours less. Find the speed going

OpenStudy (anonymous):

Did you try to figure out how to solve it?

OpenStudy (anonymous):

say the speed going is v. say it too time t to reach the destination while going. express v and t in terms of the distance.

OpenStudy (anonymous):

took* time t

OpenStudy (anonymous):

nikki, we need you to give us some feedback in order to help you.

OpenStudy (anonymous):

yeah, solving your math homework is not the point of Openstudy

OpenStudy (anonymous):

I did try...I ended up with 2 different variables and cannot solve.

OpenStudy (anonymous):

Write down what you got. We can correct you

OpenStudy (anonymous):

I got for the return trip 180=2r(t-3)

OpenStudy (anonymous):

r is the speed and t is time

OpenStudy (anonymous):

right, what is the going trip formula?

OpenStudy (anonymous):

180=r(t+3)

OpenStudy (anonymous):

I do not know the time to figure out the speed

OpenStudy (anonymous):

If im not mistaken the original should be 180 = r* (t)

OpenStudy (anonymous):

ok but how do you figure out the time to get the rate?

OpenStudy (anonymous):

why did you write 180 = r(t+3)? Does the trucker take 6 extra hours on the way to the destination?

OpenStudy (anonymous):

well simplify the second equation

OpenStudy (anonymous):

first define what t is.

OpenStudy (anonymous):

because I added three hours to the time of the return trip

OpenStudy (anonymous):

t is time

OpenStudy (anonymous):

define what t is. what do you mean by t?

OpenStudy (anonymous):

t is time taken to do what?

OpenStudy (anonymous):

the time for the total distance of 180

OpenStudy (anonymous):

time for total distance of 180? what does that mean? how can time be for distance?

OpenStudy (anonymous):

Nikki, t represents the time of the going trip

OpenStudy (anonymous):

What it the time of the going back trip?

OpenStudy (anonymous):

it takes t amount of time to go 180 miles on the going trip, and it takes t-3hours on the return trip. im confused on how to find t because i need t to find r

OpenStudy (anonymous):

I think this is a classic mechanic problem. So you can use the velocity formula (V=d/t)

OpenStudy (anonymous):

right, so it takes t amount of time, going at speed r to reach 180 miles.

OpenStudy (anonymous):

so 180 = t*r

OpenStudy (anonymous):

That is what I am using i learned it as d=rt which the same thing with only the distance in a different spot

OpenStudy (anonymous):

okay while returning, the speed is 2r, and the time take to travel back the same distance reduces by 3 hours. so 180 = 2r(t-3)

OpenStudy (anonymous):

yes i have that. but how do you find the speed from that with another variable

OpenStudy (anonymous):

so, lets call 180 = tr ------------------------- equation 1 180 = 2r(t-3) ------------------------- equation 2. from equation 1 we have t = 180/r, correct?

OpenStudy (anonymous):

yes!!! thank you i got it from here!!! much thanks

OpenStudy (anonymous):

so substitute t = 180/r in equation 2 and post what you got.

OpenStudy (anonymous):

alright, you are welcome.

OpenStudy (anonymous):

30=r